Train A and B are traveling in the same direction on parallel tracks. Train A is traveling at 100 miles per hour and train B is traveling at 120 miles per hour. Train A passes a station at 2.20 A.M. If train B oases the same station at 2:35 A.M., at what time will train B catch up to train A?
I honestly don't know where to begin solving this.

Answer by stanbon(48558)

Train A and B are traveling in the same direction on parallel tracks. Train A is traveling at 100 miles per hour and train B is traveling at 120 miles per hour. Train A passes a station at 2.20 A.M. If train B oases the same station at 2:35 A.M., at what time will train B catch up to train A?
---------------------------------
Draw the picture of the two trains moving in the same direction.
---------------
Train A DATA:
rate = 100 mph ; time = x hrs ; distance = rt = 100x miles
----------------------
Train B DATA:
rate = 120 mph ; time = x - (1/4) hr. ; distance = rt = 120x - 30 miles
==========================================
Equation:
distance A = distance B
100x = 120x - 30
20x = 30
x = 3/2 hrs
------------------------
Meeting time: 2:20 + 1 hr 30 min = 3:50
================================
Cheers,
stan H.